### v1

# Generating Arrow of Time from Three-Particle System with Classical Mechanics

Abstract
This paper investigates the mechanical origin of the arrow of time using a one-dimensional three-particle system composed of two heavy particles and a light mediator. Although the equations of motion are fully time-reversal symmetric, the energy exchange between the heavy particles through the light one exhibits irreversible relaxation. By applying the small-mass-ratio approximation ($m_{B}\ll m_{A},m_{C}$), we derive an effective equation for the energy difference $\Delta=E_{A}-E_{C}$,
$\frac{d\Delta}{dt}=-\Gamma\Delta,$
which shows exponential equilibration analogous to the thermal contact between two finite heat baths. Defining entropy as $S_{i}=\tfrac{k_{B}}{2}\ln E_{i}$, the total entropy production rate $\dot{S}_{A}+\dot{S}_{C}$ is always non-negative, indicating a macroscopic irreversibility emerging from reversible microscopic dynamics. The loss of velocity-sign information during the transformation from $(v_{A},v_{B},v_{C})$ to $(E_{A},E_{B},E_{C})$ is identified as the source of this symmetry breaking. The present model provides a minimal and transparent framework linking classical mechanics with thermodynamic irreversibility through coarse-graining.
Department of Physics and Astronomy, Faculty of Science and Technology, Tokyo University of Science, Noda City, Chiba, Japan
6222041@ed.tus.ac.jp  (long-term: xiaolinxiuping1@gmail.com)
1 Mechanical Analysis
The velocities of each particle in the three-particle system after any number of collision is known.[1] As widely known, the temperature of a system is propotion to the average of molecules’ kinetic energy. Thus, we represented particle A and C as finite thermal baths and B as thermal wall and expressed A being hotter than C at initial condition with $|v_{A0}|>|v_{C0}|$. Plus, from previous paper, the expression of velocities of each particle have a term $\overline{v_{AB}}=\frac{v_{A0}+v_{B0}}{2}$ and this makes equations more complicated. Therefore, for simplicity, we set
$v_{B0}=-v_{A0}=u>0$ (1)
so that the collisions between A and B can occur and the average of them is 0. In addition, we set the initial velocity of C as
$v_{C0}=-v<0$ (2)
so that the temperature of C increases as the collisions happen between B and C.By substituting these into previous paper’s formula, we get
$\begin{bmatrix}v_{A2k}\\ v_{B2k}\\ v_{C2k}\end{bmatrix}=\begin{bmatrix}\left(\nu^{2}v+cu\right)\frac{\sin{k\theta}}{d}-u\cos{k\theta}\\ -\left[\left(\nu^{2}v+cu\right)\frac{\sin{k\theta}}{d}-u\cos{k\theta}\right]\\ -(u+cv)\frac{\sin{k\theta}}{d}-v\cos{k\theta}\end{bmatrix}$ (3)
$\begin{bmatrix}v_{A2k+1}\\ v_{B2k+1}\\ v_{C2k+1}\end{bmatrix}=\frac{1}{M_{AB}}\begin{bmatrix}\left(m_{A}-3m_{B}\right)\left[\left(\nu^{2}v+cu\right)\frac{\sin{k\theta}}{d}-u\cos{k\theta}\right]\\ \left(m_{B}-3m_{A}\right)\left[\left(\nu^{2}v+cu\right)\frac{\sin{k\theta}}{d}-u\cos{k\theta}\right]\\ -M_{AB}\left[\left(u+cv\right)\frac{\sin{k\theta}}{d}-v\cos{k\theta}\right]\end{bmatrix}$ (4)
Now, we define below;
$\varepsilon_{A}=\frac{m_{B}}{m_{A}}$ (5)
$\varepsilon_{C}=\frac{m_{B}}{m_{C}}$ (6)
In this case, the mass of B is smaller enough than other particles, which leads that the amount of these will be <<<<1. At the same time, $c,d,\nu^{2}$ will be
$c\approx 2\varepsilon_{A}$ (7)
$d\approx\frac{\varepsilon_{A}}{\sqrt{\varepsilon_{C}}}$ (8)
$\nu^{2}\approx 1-(\varepsilon_{A}+\varepsilon_{C})$ (9)
for first order approximation. Because both $\varepsilon_{A}$ and $\varepsilon_{C}$ will be the same scale, $d$ will be closer to 0 when these amounts are small enough. Plus, the total number of collisions depends on $\theta$ therefore this also tends to 0 increasing the number of collisions sufficiently for a thermodynamic approximation. Thus,
$\frac{\sin{k\theta}}{d}\rightarrow k$ (10)
$\cos{k\theta}\rightarrow 1$ (11)
Based on these, we get
$\begin{bmatrix}v_{A2k}\\ v_{B2k}\\ v_{C2k}\end{bmatrix}\approx\begin{bmatrix}k\{[1-(\varepsilon_{A}+\varepsilon_{C})]v+2\varepsilon_{A}u\}-u\\ -k\{[1-(\varepsilon_{A}+\varepsilon_{C})]v+2\varepsilon_{A}u\}+u\\ -k(2\varepsilon_{A}v+u)-v\end{bmatrix}$ (12)
Now we calculate the ratios and differences of energy. First, the energy ratio of A and B boils down to the mass ratio because the square of velocities matches. Thus, the energy ratio is $\varepsilon_{A}$ itself and this is <<<<1 so that the energy of A is rarely transfered to B Plus, In this approximation, B is much lighter than A and C thus the relative velocities of B from A and C sill be greater. Therefore, the time taken for each cycle is approximately constant. This result shows that the mass of B is small enough and as a coclusion B can be a good channel. This is connected as thermal wall having small enough heat capacity, which let the heat from hotter bath to cooler bath be transfered almost directly based on the thermodynamic analogy. Thus, we let the time B going back and forth between A and C be a constant $\Delta t$. Now, from the law of conservation of momentum, the change of velocities through the collision of A and B under the approximation of masses are calculated as
$v_{A}^{\prime}=\frac{m_{A}-m_{B}}{M_{AB}}v_{A}+\frac{2m_{B}}{M_{AB}}v_{B}\approx v_{A}+2\varepsilon_{A}(v_{B}-v_{A})$ (13)
$v_{B}=\frac{2m_{A}}{M_{AB}}v_{A}+\frac{m_{B}-m_{A}}{M_{AB}}v_{B}\approx 2v_{A}-v_{B}+2\varepsilon_{A}(v_{A}-v_{B})$ (14)
with velocities of each particle before and after the collision $v_{A},v_{B}$ and $v_{A}^{\prime},v_{B}^{\prime}$. Therefore, the change of energy of A $\Delta E_{A}$ is
$\Delta E_{A}=\frac{1}{2}m_{A}\left(v^{\prime 2}-v^{2}\right)$
$\approx 2m_{B}v_{A}(v_{B}-v_{A})$ (15)
Similarly in the collision of B and C, the change of energy of C is, under the first order approximation,
$\Delta E_{C}\approx 2m_{B}v_{C}(2v_{A}-v_{B}-v_{C})$ (16)
From this, the difference between the change of A and C after a cycle is expressed
$\Delta E_{A}-\Delta E_{C}\approx 2m_{B}(v_{A}-v_{C})(v_{B}-v_{A}-v_{C})$ (17)
with the velocities before the cycle starts $v_{A},v_{B},v_{C}$. Therefore, the differences in the change of enregy of A and C at $k$-th cycle is
$\Delta E_{A}^{(k)}-\Delta E_{C}^{(k)}\approx 2m_{B}(v_{A2k}-v_{C2k})(v_{B2k}-v_{A2k}-v_{C2k})$
Here, we define
$\Delta_{k}\equiv E_{A}^{(k)}-E_{C}^{(k)}$ (18)
$\Lambda_{k}=\frac{\Delta E_{A}^{(k)}-\Delta E_{C}^{(k)}}{\Delta_{k}}$ (19)
and this leads this approximation;
$\Delta_{k+1}-\Delta_{k}\approx-\Lambda_{k}\Delta_{k}$ (20)
Based on this, we make
$\frac{\Delta_{k+1}-\Delta_{k}}{\Delta t}\approx-\frac{\Lambda_{k}}{\Delta t}$ (21)
being continuous and lead
$\frac{\mathrm{d}\Delta}{\mathrm{d}t}\approx-\Gamma_{k}\Delta$ (22)
where
$\Gamma_{k}\equiv\frac{\Lambda_{k}}{\Delta t}$ (23)
As $\Gamma_{k}\approx\Gamma$, we assume $\Gamma_{k}$ as a constant amount and solve the differential equation to get this;
$\Delta(t)=\Delta(0)e^{-\Gamma t}$ (24)
which means the differences of energy of two particles drops to exponential relaxation. This indicates the loss of the time reversal symmetry and implication of the flow of time going along only one direction with only the assumption of mass approximation. This apparent irreversibility arises when only the energy (not the direction of velocity) is observed. With the conversion to reverse the direction of time $t\to-t$, the velocity is also conversed similarly $v\to-v$, however, the square of it, which loses the information of sign, is unchanged between before and after the conversion. Therefore, in classical mechanics, which we directly observe the velocity, the time reversal symmetry is kept but in thermodynamics, which the energy is the target of observation, it looks lost.
2 Thermodynamic Analysis
We next assume this system as finite baths connecting. In this situation, in heat equation, the heat capacity of B can be almost neglected and the energy transfered between A and C is rarely absorbed into B. That is why this system is considered as connection of two finite bath through thermal wall. the heat equations are below;
$C_{A}\dot{T}_{A}(t)=G(T_{A}(t)-T_{C}(t))$ (25)
$C_{C}\dot{T}_{C}(t)=-G(T_{A}(t)-T_{C}(t))$ (26)
where $G$ is heat conductance and $C$ is heat capacity. The solution of these are
$T_{A}(t)=T_{eq}-\frac{C_{C}}{C_{A}+C_{C}}\Delta T(0)\exp{\left[-G\left(\frac{1}{C_{A}}+\frac{1}{C_{C}}\right)t\right]}$
$T_{C}(t)=T_{eq}+\frac{C_{A}}{C_{A}+C_{C}}\Delta T(0)\exp{\left[-G\left(\frac{1}{C_{A}}+\frac{1}{C_{C}}\right)t\right]}$ (27)
where $T_{eq}$ is the temperature at the thermal equilibrium state and $\Delta T(0)$ is the initial temperature difference. Therefore, calculating the difference of these, we get
$T_{A}(t)-T_{C}(t)=-\Delta T(0)\exp{\left[-G\left(\frac{1}{C_{A}}+\frac{1}{C_{C}}\right)t\right]}$ (28)
which has the same structure asexponential relaxation appeared in mechanical analysis.
Here, the temperatures of each particle is, because the degree of freedom is 1, connected with $E_{i}$ for particle ii(ii=A,B,C) based on law of equipartition of energy;
$E_{i}=\frac{k_{B}T_{i}}{2}$ (29)
$\therefore T_{i}=\frac{2E_{i}}{k_{B}}=\frac{m_{i}v_{i}^{2}}{k_{B}}$ (30)
This implies that we can calculate the instantaneous temperature with only the mass and velocity without depending on the average of physical quantities in case of system with few particles.
3 Statistical Mechanical Analysis
Now, we analyze this system from the perspective of statistical mechanics. According to Gibbs-Duhem’s equation,
$\frac{\partial S_{i}}{\partial E_{i}}=\frac{1}{T_{i}}$ (31)
$\therefore S_{i}=\int\frac{k_{B}}{2E_{i}}\mathrm{d}E_{i}$ (32)
$=\frac{k_{B}}{2}\ln{|E_{i}|}+C$ (33)
therefore, the ratio of entropy generation is
$\dot{S}_{i}=\frac{k_{B}}{|v_{i}|}$ (34)
B has much great $|v_{i}|$ thus the total ratio of entropy generation of A and C is
$\dot{S}_{A}+\dot{S}_{C}=k_{B}\left(\frac{1}{|v_{A}|}+\frac{1}{|v_{C}|}\right)>0$ (35)
Therefore, the law of entropy increase holds even within a purely mechanical system.
4 Discussion and Conclusion
In this study, we derived an effective thermal description from a completely mechanical three-particle system. The comparison between mechanical and thermal systems is summarized as follows:
Mechanical quantity Thermal analogue
Energies $E_{A}$, $E_{C}$ Internal energies of heat baths A and C
Particle $B$ Thermal wall (mediator of energy)
One collision cycle Infinitesimal time step dtdt
Energy difference $\Delta=E_{A}-E_{C}$ Temperature difference $\Delta T=T_{A}-T_{C}$
Coefficient $\Gamma$ Effective conductance $G(1/C_{A}+1/C_{C})$
Loss of sign $(v\rightarrow v^{2})$ Coarse-graining / irreversibility
When the mediator particle is much lighter than the other two ($m_{B}\ll m_{A},m_{C}$), the energy difference between A and C decreases monotonically. The obtained differential equation
$\frac{d\Delta}{dt}=-\Gamma\Delta$ (36)
is equivalent to the thermal relaxation equation of two finite heat baths in contact,
$\frac{d(T_{A}-T_{C})}{dt}=-G\left(\frac{1}{C_{A}}+\frac{1}{C_{C}}\right)(T_{A}-T_{C}),$ (37)
whose solution is the exponential form
$\Delta T(t)=\Delta T(0)e^{-\Gamma t}.$ (38)
This result shows that the macroscopic thermal relaxation can be described by the microscopic dynamics of elastic collisions.
Although the equations of motion are time-reversal symmetric, the transformation from $(v_{A},v_{B},v_{C})$ to $(E_{A},E_{B},E_{C})$ removes the sign of each velocity. As a result, the macroscopic description based on energy becomes irreversible. In this sense, the breaking of time-reversal symmetry arises not from the dynamics themselves but from the loss of microscopic information caused by coarse-graining.
From the definition of entropy
$S_{i}=\frac{k_{B}}{2}\ln E_{i}+\text{const},$ (39)
and the energy conservation $\dot{E}_{A}+\dot{E}_{C}\simeq 0$, the total entropy production rate is
$\dot{S}_{A}+\dot{S}_{C}=k_{B}\left(\frac{1}{|v_{A}|}+\frac{1}{|v_{C}|}\right)>0$ (40)
Therefore, the entropy of the system always increases, even though the underlying dynamics are reversible.
In conclusion, the present model demonstrates that the arrow of time can appear in a deterministic mechanical system when information about velocity direction is lost. This provides a simple and quantitative example that connects classical mechanics with thermodynamics through coarse-graining. Further work may extend this model to multi-particle or quantum systems to explore collective and microscopic origins of irreversibility.This work thus bridges microscopic reversibility and macroscopic irreversibility within a single deterministic framework.